OpenStudy (anonymous):

What is the perimeter of a square with a diagonal of 3√2?

OpenStudy (anonymous):

3 , 12, or 18

OpenStudy (anonymous):

12

OpenStudy (anonymous):

thank u!

OpenStudy (saifoo.khan):

12

OpenStudy (anonymous):

If you use Pythagorean Theorum the hypotenuse squared is going to equal the squares of the other two sides (which are equal). So, (3√2)^2 = x^2 + x^2 9(2) = 2x^2 18 = 2x^2 9 = x^2 3 = x, so each side is 3 units long. A square has four sides, so the perimeter is 3 x 4 or 12 units in length.